An investment project provides cash inflows of $620 per year for 9 years.

a. What is the project payback period if the initial cost is $1,860?

b. What is the project payback period if the initial cost is $3,534?

c. What is the project payback period if the initial cost is $6,200?
Answer #1

payback period=initial cost/annual cash flows

1.payback period=1860/620=3 years.

2.payback period=3534/620=5.7 years

3.payback period=6200/620=Never